OpenStudy (anonymous):

find the limit of this please \[\lim_{x \rightarrow -infinity}(5x+12)/\sqrt{9x^2+x-3}\]

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}\frac{5x+12}{\sqrt{9x^2+x-3}} \cdot \frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}}\] now recall \[\sqrt{x^2}=|x|=x \text{ if } x<0 \] x<0 since we are approaching negative infinity so we have \[\lim_{x \rightarrow -\infty}\frac{5x+12}{\sqrt{9x^2+x-3}} \cdot \frac{\frac{1}{-x}}{ \frac{1}{\sqrt{x^2}}}\]

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty} \frac{-5+\frac{-12}{x}}{\sqrt{9+\frac{1}{x}-\frac{3}{x^2}}}\]

myininaya (myininaya):

\[\frac{-5 +0}{\sqrt{9+0-0}}\]

myininaya (myininaya):

oops small type-0 |x|=-x if x<0

myininaya (myininaya):

everything else is good

OpenStudy (anonymous):

is the final answer -5/sqrt 9

myininaya (myininaya):

the sqrt(9)=3

OpenStudy (anonymous):

ok i see thanks again myininaya

myininaya (myininaya):

np